OpenStudy (anonymous):

A tennis court is 36 feet wide. Its length is 6 feet more than twice the width. What is the area in square feet of a tennis court?

OpenStudy (anonymous):

x=15

OpenStudy (anonymous):

is that the area?

OpenStudy (anonymous):

540 sq. ft.

OpenStudy (anonymous):

Length = Width * 2 + 6 Length = 36 * 2 + 6 Length = 78 Area = Width * Length Area = 36 * 78 Area = 2808

OpenStudy (anonymous):

i thought length is given,actually width is given.